By analyzing the gametes from a trihybrid fruit fl
Answer #1

Parental: ABD

abd

Single cross over between genes D and B AbD

aBd

Double cross over ABd

abD

Single cross over between genes A and D Abd

aBD

Distance between genes A and D = 37+44+3+6

----------------- X100 = 9cM

1000

Distance between genes D and B = 43+48+3+6

---------------- 100 = 10cM

1000

interference = 1 -( observed DCO/expected DCO)

   1-(3+6)/(0.1)(0.09) (1000)

   1-9/9=0

complete chromosomal interference observed

I